OpenStudy (anonymous):

Solve the system of linear equations by the substitution method. please show. =) x = 3y + 30 3x + 6y = 0

OpenStudy (anonymous):

3(3y+30)+6y=0 9y+90+6y=0 15y=-90 y=-6 x = 3y + 30 x=3(-6)+30 x=-12

OpenStudy (anonymous):

You owe $976.34 on a credit card that has an interest rate of 10.75% APR. You pay $100.00 at the end of each month. You place the $100.00 in a savings account that earns a 2.75% APR. What is the difference in interest between savings earned and credit card interest paid?

OpenStudy (anonymous):

sorry x=12 NOT -12

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Your welcome :)